# Transfer and routing of Gaussian states through quantum complex networks with and without community structure

### v1

@@ Abstract
The goal in quantum state transfer is to avoid the need to physically transport carriers of quantum information. This is achieved by using a suitably engineered Hamiltonian that induces the transfer of the state of one subsystem to another. A less known generalization of state transfer considers multiple systems such that any pair can exchange quantum information and transfers can take place at any time, starting and stopping independently. This is sometimes called routing of quantum states. State transfer in particular has received a great deal of attention, however the vast majority of results in both state transfer and routing concern qubits transferred in a network of restricted structure. Here we consider routing of single-mode Gaussian states and entanglement through complex networks of quantum harmonic oscillators. We compare a protocol where the transfer is completed in a single step but the effective Hamiltonian only approximately transfers the state with one where the transfer can in principle be perfect but the transfer is done in two steps, and also illustrate the state-dependency of the transfer fidelity. We find that even in a random and homogeneous network, the transfer fidelity still depends on the degree of the nodes for any link density, and that in both random and complex networks it is the community structure that controls the appearance of higher frequency normal modes useful for transfer. Finally, we find that networks of sufficient complexity may have superior routing performance over superficially similar random networks. Our results pave the way for further exploration of the role of community structure in state transfer and related tasks.
preprint: APS/123-QED
I Introduction
@@
Quantum state transfer is the task of transferring a state from some initial system to a target system through Hamiltonian dynamics of a network of many interacting systems. This could be used to realize high fidelity channels connecting different quantum processors, perhaps eventually leading to all solid state quantum information processing on a chip [1, 2, 3]. State transfer is also considered from a more formal point of view as a fundamental problem in mathematical physics [4, 5]. Sometimes networks with special symmetries that can naturally transfer the state are sought. This can potentially lead to a fast transfer with minimal control requirements, though heavy restrictions are imposed on the network [6, 7, 4, 8]. Alternatively, the network may be given and one tries to achieve transfer assuming some limited control over a small part of it, typically by engineering a situation where the sender and receiver effectively interact with only a single normal mode of the rest of the network which then acts as a communication channel [9, 10, 11, 12]. In this approach the standard choice is the normal mode corresponding to the smallest eigenvalue of the network, since in the case that the Hamiltonian is proportional to the Laplace matrix of the network, the mode has by construction uniform overlap with all the nodes [13], and for example in Erdős-Rényi networks it is well isolated from the other normal modes due to a large spectral gap [14]. Both approaches have been used in fermionic systems, often to transfer single qubits [6, 15, 16], and in bosonic systems, notably to transfer of entanglement [17, 9, 18], number and coherent states [8] or excitations [19]. In the continuous variable case state transfer is of particular interest when distances are short since unlike with qubits, perfect teleportation is impossible [20, 21]. In general, it can be thought of as a two-user protocol where the users are the sender and the receiver of the state to be transferred. Logically, this corresponds to a two-port bus connecting the sender and the receiver. In particular, even if the users can be freely chosen, only a single transfer takes place at any given time.
@@
We address these questions by considering transfer and routing of Gaussian states and entanglement in both random and empirical bosonic networks. We compare and contrast two protocols, observing also how a resource such as squeezing or entanglement may be transferred quite well even if phase differences cause the fidelity to be relatively low. We show how even in Erdős-Rényi networks—considered largely featureless as networks—the external users are not equal. Specifically the degree of the node the user couples to correlates with the fidelity independently of the link density. This makes an abundance of high degree nodes beneficial overall, and therefore also the degree distribution relevant to the transfer fidelity. We then reveal the crucial role of community structure in facilitating routing in complex networks, a network property that has formerly remained mostly unexplored also in state transfer. Considering a social network with a known community structure, we further identify which nodes allow high fidelity transfer over given normal modes to reveal the finer structure of the available communication channels. This also finds a finer community structure with high modularity as a by-product. Finally, we introduce tentative measures for routing capacity and find that a sufficiently complex network may outperform a random network with superficially similar features.
@@
The rest of this work is structured as follows. In Sec. II we introduce the model and the scheme for using it to route single-mode Gaussian states and entanglement. We also compare the transfer performance and speed of the one and two step transfer protocols, and how the former depends on the chosen state and the used figure of merit. In Sec. III we consider random networks both without and with a community structure, revealing its role in the routing capabilities of the network. Similar results are also found for some empirical networks, which we consider in Sec. IV. Here we also benchmark the empirical networks to their randomized counterparts. We conclude in Sec. V where we also discuss new questions our results have raised.
II Quantum state routing
@@
Figure 1: The routing problem and some different approaches: $a)$ a ring, $b)$ a random network with homogeneous link density and $c)$ a network with a community structure. The white circles are users that can act as both senders and receivers of quantum information. Enclosed in the big circle is a central system that should facilitate routing such that its Hamiltonian (consisting of subsystems and interaction terms indicated by the blue circles and lines, respectively) can remain fixed and there is no need to control its state. Only local control of the users and their couplings to the central system, indicated by the black lines, should suffice to facilitate independent and possibly overlapping transfers between any pairs of users.
II.1 The model
@@
where $\rho_{1}$ and $\rho_{2}$ are density operators of the initial and transferred state. We will also use efficiency, which is the ratio between the transferred and initial amount of a specific resource such as squeezing or entanglement. For single mode Gaussian states considered here the fidelity can be easily calculated directly from the covariance matrices and first moment vectors.
II.2 State transfer over the normal modes
@@
Figure 3: Comparison of transfer performance for different states and transfer speeds for the single step protocol in the ideal case of only a single normal mode. Transfer time increases linearly with the integer $c_{3}\geq 1$ but performance improves. Efficiency is shown in cases $a)$, $c)$ and $e)$ and fidelity in the others. The states are a coherent state with magnitude of displacement $|\alpha|$ in $a),b)$ and a squeezed vacuum with squeezing parameter $r$ in $c),d)$. A twin beam state with two-mode squeezing parameter $s$ is also considered such that only one of the systems couples to the chain; shown are the fraction of transferred entanglement quantified by logarithmic negativity $E_{\mathcal{N}}$ in $e)$ and fidelity with a twin beam state with the same $s$ but orthogonal phase in $f)$.
@@
We consider two inequivalent ways to transfer the state, accomplished by judiciously tuning the effective Hamiltonian and interaction times as explained in Appendix B. In the two step protocol only the sender couples to the network first, and swaps states with the normal mode. Next only the receiver couples to the network and also swaps states with the normal mode. The state is transferred in principle perfectly to the receiver, whereas the original states of the normal mode and the receiver end up in the sender and the normal mode, respectively. In the single step protocol both the sender and the receiver couple to the network simultaneously and for the same duration such that their states are swapped while the normal mode state remains invariant. Both protocols are parameterized by a positive integer such that higher values decrease the coupling strengths but increase transfer times as explained in Appendix B, namely $c_{2}$ and $c_{3}$ according to whether the transfer involves two coupled systems at a time or all three. On the one hand the single step protocol only approximately achieves the transfer even in the ideal case where the other normal modes are not present, but on the other hand can also be used for two-way transfer as the external systems can both simultaneously send and receive quantum information. We remark that achieving two-way transfer with the first protocol would require an additional third step where the sender again couples to the network to once more swap states with the normal mode. For now, we focus on the ideal case where the effect of non-resonant normal modes vanish, addressing the role of the network structure in later Sections.
@@
The two protocols are compared in Fig. 2. The state to be transferred is $\left|r=0.5,\varphi=0\right>$ and the figure of merit is the efficiency of squeezing transfer, i.e. $r(t)/r(0)$ where $r(0)=0.5$. The differences become apparent during the evolution, as does the approximate nature of the single step protocol. In particular, although the transfer time for it should be such that $t\omega_{0}=3\pi$ as suggested by Eq. (10), it can be observed that in practice the efficiency reaches its maximum value a bit later. This suggests a simple optimization strategy for the transfer time for this protocol which will be used later in Sec. III. Although the single step protocol is in principle slightly faster we show in Appendix C that asymptotically the transfer times are identical whereas the coupling strength required by the two step protocol is stronger by a factor of $\sqrt{2}$.
@@
III Erdős-Rényi and modular random networks
In this section we present the routing capabilities of Erdős-Rényi (ER) and modular random networks when the single step transfer protocol is used, finding that the degree of connection nodes matters to transfer fidelity, and that the community structure of modular networks allows transfer over frequencies other than the-center-of mass mode. Although in the ideal case the two step protocol allows perfect state transfer, it is more vulnerable to the presence of additional normal modes due to the stronger coupling it requires, as shown in Appendix C. In the general case, the performance difference is in favor of the single step protocol.
III.1 Slowest normal mode in Erdős-Rényi random networks
@@
The oscillators have unit mass and frequency, and the coupling strength between network oscillators is 1. The external oscillators are coupled to their respective network nodes with strength $k=g/\sqrt{n}$, where $g$ corresponds to Eq. (10) with $c_{3}=7$. This equates to an effective coupling strength of $g$ to the slowest normal mode. Similarly for the transfer time the value from the ideal case $t_{ideal}$ was used. However, due to the idealized assumptions in the derivation of the transfer time, the maximum fidelity is not reached exactly at the calculated time but slightly after due to an unpredictable phase difference. In addition, the difference between the times varies between realizations. The system was thus simulated for time window $[t_{ideal},t_{ideal}+4]$ to capture a full period of the oscillator and thus at least one instant where the phases match. The maximum fidelity within this window was taken as the transfer fidelity.
In order to quantify the importance of both connection nodes’ degree, we consider the transfer fidelity as a function of their average degree, $k_{avg}=(k_{S^{\prime}}+k_{R^{\prime}})/2$. In each ensemble all pairs with the same $k_{avg}$ were grouped together. The mean of the transfer fidelities achieved between such pairs are presented in Fig. 4 for each ensemble. We observe that a higher average degree of the connection nodes translates to better transfer fidelity, while the degrees of the other nodes in the network do not affect the fidelity. The degree of a node increases its effective frequency, which in turn weakens the coupling to lower frequency normal modes. Higher degree nodes thus couple more weakly to the normal modes closer to the center-of-mass mode, leading to smaller loss of information to the other normal modes.
@@
Figure 5: Transfer fidelity of a squeezed vacuum, in each ensemble averaged over pairs with the same degrees on both senders and receivers connection nodes $S^{\prime}$ and $R^{\prime}$. Ensembles with $p=0.2$, $0.5$ and $0.8$ are presented. For lower degree nodes a bottleneck effect is visible: the connection node with lower degree limits the transfer fidelity for the pair. Due to the Poissonian degree distribution of ER-networks, pairs in the middle of each ensembles mass are more common than those at the perimeter.
Note that the colour scale is again composed of two linear ranges above and below $F=0.95$.
III.2 Normal modes in modular random networks
@@
A network with community structure may allow transfer over normal modes other than the center of mass mode. In general, a network with $M$ communities has, including the center of mass mode, $M$ normal modes which are well separated from the rest and to which multiple nodes couple with significant strength [36]. The coupling strengths to these normal modes depend strongly on the community the node is in. State transfer is then possible within a community over those modes to which the nodes couple strongly, and between communities if the nodes in both communities couple strongly to the same normal mode. If the communities have no good fast normal modes in common, they may in any case communicate over the slowest normal mode.
@@
To generate modular random networks we used a stochastic block model (SBM) [37]. In effect the model constructs $M$ communities of ER-networks, and randomly adds edges between the communities. We present here results for a 40-node network of four equally sized communities, with edge probability $p_{\text{w}}=0.75$ within each community and $p_{\text{b}}=0.025$ between communities. A single example realization is shown in Fig. 6, with the mean coupling strengths of the nodes in each community to the four lowest normal modes shown in the inset. As noted earlier, by construction all nodes couple to the center of mass mode with equal strength. For the other modes the coupling strength is strongly dependent on the community, and at least two communities have a significant coupling to each normal mode. We observe that this behaviour holds for most networks with the chosen parameters.
@@
The oscillators have same properties as in Sec. III.1, and the external oscillators’ effective coupling strength is according to Eq. (10) with $c_{3}=50$. The large value for $c_{3}$, corresponding to a very weak coupling, was chosen to highlight the difference between good and bad normal modes; with a strong coupling, even good normal modes could allow only low fidelity transfer. The state to be transferred is again the squeezed vacuum $\left|r=1,\varphi=0\right>$.
@@
Fig. 7 shows the mean transfer fidelities over the lowest four normal modes for an ensemble of 200 random modular networks. As expected, over the center of mass mode transfer between any pair in the network reaches high fidelity. Over the higher normal modes there is in general a subgraph of only two communities in which high fidelity transfer is possible, while fidelities are low between pairs where one or both participants do not couple to this subgraph. In particular, high fidelity in the ‘Top two’ row indicates the possibility to communicate both within and between different communities. The communities which reach high fidelities may be identified as the communities which couple strongly to the normal mode.
@@
IV Empirical networks
IV.1 Role of community structure
Figure 8: Mean transfer fidelity for a squeezed vacuum $\left|r=1,\varphi=0\right>$ between the original communities of the Karate club network. The fidelities are presented for the five slowest normal modes.
Figure 9: By grouping nodes according to transfer fidelities, we may find further community structure in the karate club network. This partition, indicated by the different colours, splits both original communities into two: nodes marked with a triangle correspond to original community 1 and nodes with a circle to community 2.

### v2

@@ Abstract
The goal in quantum state transfer is to avoid the need to physically transport carriers of quantum information. This is achieved by using a suitably engineered Hamiltonian that induces the transfer of the state of one subsystem to another. A less known generalization of state transfer considers multiple systems such that any pair can exchange quantum information and transfers can take place at any time, starting and stopping independently. This is sometimes called routing of quantum states. State transfer in particular has received a great deal of attention, however the vast majority of results in both state transfer and routing concern qubits transferred in a network of restricted structure. Here we consider routing of single-mode Gaussian states and entanglement through complex networks of quantum harmonic oscillators. We compare a protocol where the transfer is completed in a single step but the effective Hamiltonian only approximately transfers the state with one where the transfer can in principle be perfect but the transfer is done in two steps, and also illustrate the state-dependency of the transfer fidelity with paradigmatic Gaussian states as well as number states. We find that even in a random and homogeneous network, the transfer fidelity still depends on the degree of the nodes for any link density, and that in both random and complex networks it is the community structure that controls the appearance of higher frequency normal modes useful for transfer. Finally, we find that networks of sufficient complexity may have superior routing performance over superficially similar random networks. Our results pave the way for further exploration of the role of community structure in state transfer and related tasks.
1 Introduction
@@
Quantum state transfer is the task of transferring a state from some initial system to a target system through Hamiltonian dynamics of a network of many interacting systems. This could be used to realize high fidelity channels connecting different quantum processors, perhaps eventually leading to all solid state quantum information processing on a chip [1, 2, 3]. State transfer is also considered from a more formal point of view as a fundamental problem in mathematical physics [4, 5]. Sometimes networks with special symmetries that can naturally transfer the state are sought. This can potentially lead to a fast transfer with minimal control requirements, though heavy restrictions are imposed on the network [6, 7, 4, 8]. Alternatively, the network may be given and one tries to achieve transfer assuming some limited control over a small part of it, typically by engineering a situation where the sender and receiver effectively interact with only a single normal mode of the rest of the network which then acts as a communication channel [9, 10, 11, 12]. In this approach the standard choice is the normal mode corresponding to the smallest eigenvalue of the network, since in the case that the Hamiltonian is proportional to the Laplace matrix of the network, the mode has by construction uniform overlap with all the nodes [13], and for example in Erdős-Rényi networks it is well isolated from the other normal modes due to a large spectral gap [14] which has been shown to be relevant to state transfer [9]. Both approaches have been used in fermionic systems, often to transfer single qubits [6, 15, 16], and in bosonic systems, notably to transfer of entanglement [17, 9, 18], number and coherent states [8] or excitations [19]. In the continuous variable case state transfer is of particular interest when distances are short since unlike with qubits, perfect teleportation is impossible [20, 21]. In general, it can be thought of as a two-user protocol where the users are the sender and the receiver of the state to be transferred. Logically, this corresponds to a two-port bus connecting the sender and the receiver. In particular, even if the users can be freely chosen, only a single transfer takes place at any given time.
@@
We address these questions by considering transfer and routing of Gaussian states and entanglement in both random and empirical bosonic networks. We compare and contrast two protocols, observing also how a resource such as squeezing or entanglement may be transferred quite well even if phase differences cause the fidelity to be relatively low. We also confirm the applicability of our scheme to non-Gaussian states by demonstrating high transfer fidelities for number states. We show how even in Erdős-Rényi networks—considered largely featureless as networks—the external users are not equal. Specifically the degree of the node the user couples to correlates with the fidelity independently of the link density. This makes an abundance of high degree nodes beneficial overall, and therefore also the degree distribution relevant to the transfer fidelity.
We then reveal the crucial role of community structure in facilitating routing in complex networks, a network property that has formerly remained mostly unexplored also in state transfer. Community structure, an important property in many real networks, is characterized by a non-uniform link density [33]. It breaks the network into densely connected subnetworks—the communities—with substantially fewer links connecting nodes belonging to different communities. This is exemplified in Fig. 1 c) where the communities are circled. Since no formal definition exists, there is no canonical way to find the communities of a network. How well a given division does may be quantified through different measures such as modularity [34, 35]. Considering a social network with a known community structure, we further identify which nodes allow high fidelity transfer over given normal modes to reveal the finer structure of the available communication channels. This also finds a finer community structure with high modularity as a by-product. Finally, we introduce tentative measures for routing capacity and find that a sufficiently complex network may outperform a random network with superficially similar features.
@@
The rest of this work is structured as follows. In Sec. 2 we introduce the model and the scheme for using it to route single-mode Gaussian states and entanglement. We also compare the transfer performance and speed of the one and two step transfer protocols, and how the former depends on the chosen state and the used figure of merit. In Sec. 3 we consider random networks both without and with a community structure, revealing its role in the routing capabilities of the network. Similar results are also found for some empirical networks, which we consider in Sec. 4. Here we also benchmark the empirical networks to their randomized counterparts. We conclude in Sec. 5 where we also discuss new questions our results have raised.
2 Quantum state routing
[…]
[trim=0cm 11cm 7.5cm 0cm,clip=true,width=0.48]figures/Figure1purple.pdf
@@ 2 Quantum state routing
Figure 1: The routing problem and some different approaches: $a)$ a ring, $b)$ a random network with homogeneous link density and $c)$ a network with a community structure, with the three communities circled with dashed line. The white circles are users that can act as both senders and receivers of quantum information. Enclosed in the big circle is a central system that should facilitate routing such that its Hamiltonian (consisting of subsystems and interaction terms indicated by the blue circles and lines, respectively) can remain fixed and there is no need to control its state. Only local control of the users and their couplings to the central system, indicated by the black lines, should suffice to facilitate independent and possibly overlapping transfers between any pairs of users.
2.1 The model
@@
where $\rho_{1}$ and $\rho_{2}$ are density operators of the initial and transferred state. We will also use efficiency,
$\mathrm{eff}_{X}(t)=\frac{X_{R}(t)}{X_{S}(0)}$ (3)
which is the ratio of a specific resource $X$ that was transferred to the receiver $R$ at time $t$ to the initial amount at the sender $S$. The resource $X$ can be for example squeezing or entanglement. For single mode Gaussian states considered here the fidelity can be easily calculated directly from the covariance matrices and first moment vectors.
2.2 State transfer over the normal modes
[…]
[trim=0cm 0cm 0.0cm 0cm,clip=true,width=0.48]figures/idealtransfer_revised.pdf
@@ 2.2 State transfer over the normal modes
[…]
[trim=0cm 0cm 0.0cm 0cm,clip=true,width=0.48]figures/fidelity_scaling_new.pdf
Figure 3: Comparison of transfer performance for different states and transfer speeds for the single step protocol in the ideal case of only a single normal mode. Transfer time increases linearly with the integer $c_{3}\geq 1$ but performance improves. Efficiency is shown in panels $a)$, $c)$, $e)$ and $g)$ and fidelity in the others. Panels $a)$ $f)$ consider transfers of single mode states: a coherent state with magnitude of displacement $|\alpha|$ in $a)$, $b)$, a squeezed vacuum with squeezing parameter $r$ in $c)$, $d)$, and a number state $|n\rangle$ in $e)$, $f)$. Notably for the number state, which is non-Gaussian, the behavior is similar to the Gaussian states. A twin beam state with two-mode squeezing parameter $s$ is also considered in $g)$, $h)$, such that only one of the systems couples to the chain; shown are the fraction of transferred entanglement quantified by logarithmic negativity $E_{\mathcal{N}}$ in $g)$ and fidelity with a twin beam state with the same $s$ but orthogonal phase in $h)$.
@@
We consider two inequivalent ways to transfer the state, accomplished by judiciously tuning the effective Hamiltonian and interaction times as explained in App. B. Detection or state tomography is not considered as it has no effect on the transfer fidelity. In the two step protocol only the sender couples to the network first, and swaps states with the normal mode. Next only the receiver couples to the network and also swaps states with the normal mode. The state is transferred in principle perfectly to the receiver, whereas the original states of the normal mode and the receiver end up in the sender and the normal mode, respectively. In the single step protocol both the sender and the receiver couple to the network simultaneously and for the same duration such that their states are swapped while the normal mode state remains invariant. Both protocols are parameterized by a positive integer such that higher values decrease the coupling strengths but increase transfer times as explained in App. B, namely $c_{2}$ and $c_{3}$ according to whether the transfer involves two coupled systems at a time or all three. On the one hand the single step protocol only approximately achieves the transfer even in the ideal case where the other normal modes are not present, but on the other hand can also be used for two-way transfer as the external systems can both simultaneously send and receive quantum information. We remark that achieving two-way transfer with the first protocol would require an additional third step where the sender again couples to the network to once more swap states with the normal mode. For now, we focus on the ideal case where the effect of non-resonant normal modes vanish, addressing the role of the network structure in later Sections.
@@
The two protocols are compared in Fig. 2. The state to be transferred is $\left|r=0.5,\varphi=0\right>$ and the figure of merit is the efficiency of squeezing transfer $\mathrm{eff}_{r}(t)$. The differences become apparent during the evolution, as does the approximate nature of the single step protocol. In particular, although the transfer time for it should be such that $t\omega_{0}=3\pi$ as suggested by Eq. (11), it can be observed that in practice the efficiency reaches its maximum value a bit later. This suggests a simple optimization strategy for the transfer time for this protocol which will be used later in Sec. 3. Although the single step protocol is in principle slightly faster we show in App. C that asymptotically the transfer times are identical whereas the coupling strength required by the two step protocol is stronger by a factor of $\sqrt{2}$.
@@
We also consider transfer of a number state in Fig. 3e), f) as an example of a non-Gaussian state. As noted in App. B, in the two step protocol the Gaussian nature of the transferred state is irrelevant to the success of the transfer; this seems to extend to the one step protocol as seen from the the transfer efficiencies’ behavior. However, compared to the Gaussian states a high fidelity is easier to achieve for the number states due to them being phase invariant. The numerical simulations for the number states are performed differently from the Gaussian states and the details regarding the simulations are included in App. D.
@@
3 Erdős-Rényi and modular random networks
In this section we present the routing capabilities of Erdős-Rényi (ER) and modular random networks when the single step transfer protocol is used, finding that the degree of connection nodes matters to transfer fidelity, and that the community structure of modular networks allows transfer over frequencies other than the-center-of mass mode. Although in the ideal case the two step protocol allows perfect state transfer, it is more vulnerable to the presence of additional normal modes due to the stronger coupling it requires, as shown in App. C. In the general case, the performance difference is in favor of the single step protocol.
3.1 Slowest normal mode in Erdős-Rényi random networks
[…]
[trim=0cm 0cm 0.0cm 0cm,clip=true,width=0.48]figures/er_fidelity_p_avgdegree.pdf
@@ 3.1 Slowest normal mode in Erdős-Rényi random networks
The oscillators have unit mass and frequency, and the coupling strength between network oscillators is 1. The external oscillators are coupled to their respective network nodes with strength $g^{\prime}=g/\sqrt{N}$, where $g$ corresponds to Eq. (11) with $c_{3}=7$ and $N=30$ is the number of oscillators. This equates to an effective coupling strength of $g$ to the slowest normal mode. Similarly for the transfer time the value from the ideal case $t_{ideal}$ was used. However, due to the idealized assumptions in the derivation of the transfer time, the maximum fidelity is not reached exactly at the calculated time but slightly after due to an unpredictable phase difference. In addition, the difference between the times varies between realizations. The system was thus simulated for time window $[t_{ideal},t_{ideal}+4]$ to capture a full period of the oscillator and thus at least one instant where the phases match. The maximum fidelity within this window was taken as the transfer fidelity.
In order to quantify the importance of both connection nodes’ degree, we consider the transfer fidelity as a function of their average degree, $k_{avg}=(k_{S^{\prime}}+k_{R^{\prime}})/2$. In each ensemble all pairs with the same $k_{avg}$ were grouped together. The mean of the transfer fidelities achieved between such pairs are presented in Fig. 4 for each ensemble. We observe that a higher average degree of the connection nodes translates to better transfer fidelity, while the overall degree distribution, and consequently the spectral gap, of the network does not affect the fidelity. The degree of a node increases its effective frequency, which in turn weakens the coupling to lower frequency normal modes. Higher degree nodes thus couple more weakly to the normal modes closer to the center-of-mass mode, leading to smaller loss of information to the other normal modes. The importance of the spectral gap is also highlighted. In networks with small spectral gaps—such as rings and paths—it must be accounted for with a weak coupling [9], but when the gap is above some threshold, its size does not affect the fidelity.
@@
[…]
[trim=0cm 0cm 0.0cm 0cm,clip=true,width=0.48]figures/er_fidelity_degrees.pdf
@@
Figure 5: Transfer fidelity of a squeezed vacuum, in each ensemble averaged over pairs with the same degrees on both senders and receivers connection nodes $S^{\prime}$ and $R^{\prime}$. Ensembles with $p=0.2$, $0.5$ and $0.8$ are presented. For lower degree nodes a bottleneck effect is visible: the connection node with lower degree limits the transfer fidelity for the pair. Due to the Poissonian degree distribution of ER-networks, pairs in the middle of each ensembles mass are more common than those at the perimeter. Note that the colour scale is again composed of two linear ranges above and below $F=0.95$.
3.2 Normal modes in modular random networks
@@
Here we consider modular networks, or networks with community structure. These networks are characterized by high link density within some groups of nodes, and a lower link density between the groups. Such networks may allow transfer over normal modes other than the center of mass mode. In general, a network with $M$ communities has, including the center of mass mode, $M$ normal modes which are well separated from the rest and to which multiple nodes couple with significant strength [39]. The coupling strengths to these normal modes depend strongly on the community the node is in. State transfer is then possible within a community over those modes to which the nodes couple strongly, and between communities if the nodes in both communities couple strongly to the same normal mode. If the communities have no good fast normal modes in common, they may in any case communicate over the slowest normal mode.
@@
To generate modular random networks we used a stochastic block model (SBM) [40]. In effect the model constructs $M$ communities of ER-networks, and randomly adds edges between the communities. We present here results for a 40-node network of four equally sized communities, with edge probability $p_{\text{w}}=0.75$ within each community and $p_{\text{b}}=0.025$ between communities. These probabilities were chosen to produce a pronounced community structure, with nodes having a high degree from links internal to their community in line with the results of the previous section. A single example realization is shown in Fig. 6, with the mean coupling strengths of the nodes in each community to the four lowest normal modes shown in the inset. As noted earlier, by construction all nodes couple to the center of mass mode with equal strength. For the other modes the coupling strength is strongly dependent on the community, and at least two communities have a significant coupling to each normal mode. We observe that this behaviour holds for most networks with the chosen parameters.
@@
The oscillators have same properties as in Sec. 3.1, and the external oscillators’ effective coupling strength is according to Eq. (11) with $c_{3}=50$. The large value for $c_{3}$, corresponding to a very weak coupling, was chosen to highlight the difference between good and bad normal modes; with a strong coupling, even good normal modes could allow only low fidelity transfer. The state to be transferred is again the squeezed vacuum $\left|r=1,\varphi=0\right>$.
@@
Figure 7 shows the mean transfer fidelities over the lowest four normal modes for an ensemble of 200 random modular networks. As expected, over the center of mass mode transfer between any pair in the network reaches high fidelity. Over the higher normal modes there is in general a subgraph of only two communities in which high fidelity transfer is possible, while fidelities are low between pairs where one or both participants do not couple to this subgraph. In particular, high fidelity in the ‘Top two’ row indicates the possibility to communicate both within and between different communities. The communities which reach high fidelities may be identified as the communities which couple strongly to the normal mode.
[…]
[trim=0cm 0cm 0.0cm 0cm,clip=true,width=0.48]figures/modular_example_inset.pdf
@@
[…]
[trim=0cm 0cm 0.0cm 0cm,clip=true,width=0.48]figures/modular_fidelity_hm.pdf
@@
4 Empirical networks
4.1 Role of community structure
[…]
[trim=0cm 0cm 0.0cm 0cm,clip=true,width=0.48]figures/karate_fids.pdf
Figure 8: Mean transfer fidelity for a squeezed vacuum $\left|r=1,\varphi=0\right>$ between the original communities of the Karate club network. The fidelities are presented for the five slowest normal modes, from $\omega_{0}$ to $\omega_{4}$.
[…]
[trim=0cm 0cm 0.0cm 0cm,clip=true,width=0.48]figures/karate_communities_graph.pdf
Figure 9: In the Karate club network, originally consisting of two communities marked here with triangles and circles, further community structure can be found by considering the transfer fidelities. Here we identify one possible partition, indicated by the different colors and letters a - d. Fidelities for this partition are shown in Fig. 10.
[…]
[trim=0cm 0cm 0.0cm 0cm,clip=true,width=0.48]figures/karate_fids_four_comms.pdf